### v1

# Quantum hydrodynamics of a polariton fluid: pure energy relaxation terms

Abstract
Cavity polaritons, hybrid half-light half-matter excitations in quantum microcavities in the strong-coupling regime demonstrate clear signatures of quantum collective behavior, such as analogues of BEC and superfluidity at surprisingly high temperatures. The analysis of the formation of these states demands an account of the relaxation processes in the system. Although there are well-established approaches for the description of some of them, such as finite lifetime polariton, an external optical pump, and coupling with an incoherent excitonic reservoir, the treatment of pure energy relaxation in a polariton fluid still remains a puzzle. Here, based on the quantum hydrodynamics approach, we derive the corresponding equations where the energy relaxation term appears naturally. We analyze in detail how it affects the dynamics of polariton droplets and the dispersion of elementary excitations of a uniform polariton condensate. Although we focus on the case of cavity polaritons, our approach can be applied to other cases of bosonic condensates, where the processes of energy relaxation play an important role.
preprint: APS/123-QED
The physics of quantum fluids represents a major part of modern condensed matter and atomic physics. In the low-temperature limit the ensembles of identical quantum particles can form macroscopically coherent states corresponding to Bose-Einstein condensates (BECs) and demonstrate the remarkable property of superfluidity. This phenomenon is well studied in the domain of cold atomic gases, where the dynamics of condensate droplets can be modeled by the Gross-Pitaevskii equation for the macroscopic wavefunction (order parameter) of the condensate [1]:
$i\hbar\frac{\partial\psi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi+U\psi+g|\psi|^{2}\psi.$ (1)
The first term on the right-hand side corresponds to a kinetic energy of the condensed particles, the second term describes an interaction with an external potential corresponding to an optical or magnetic trap, and the last non-linear term corresponds to interatomic interactions treated within the framework of the mean-field and s-wave scattering approximations. Gross-Pitaevskii equation gives a perfect description of the evolution of a conservative system with both the number of particles $N=\int|\psi(\mathbf{r},t)|^{2}d^{2}\mathbf{r}$ and the total energy:
$H=\int\left[\frac{\hbar^{2}}{2m}|\nabla\psi|^{2}+U|\psi|^{2}+\frac{g}{2}|\psi|^{4}\right]d^{2}\mathbf{r}.$ (2)
The characteristic temperatures of the BEC phase transition for cold atomic gases lie in the nanokelvin diapason, which is related to the large mass of condensing atoms, such as isotopes of sodium and rubidium. This motivated the search of analogues of BECs in condensed matter systems, where effective masses of various bosonic quasiparticles are several orders of magnitude smaller, and one can expect formation of quantum collective states at much higher temperatures.
One of the attractive possibilities is the use of ensembles of exciton polaritons (also known as cavity polaritons), which appear when a strong coupling regime between excitonic and photonic modes is realized in an optical microresonator [2, 3]. Being composite half-light-half-matter particles, polaritons inherit an extremely small effective mass (about $10^{-5}$ of the mass of free electrons) and large coherence length (in the mm scale) [4] from their photonic component. However, the presence of an excitonic component results in efficient polariton-polariton interactions, which lead to an extremely strong nonlinear optical response. The polariton BEC and superfluidity were experimentally observed at surprisingly high temperatures under both optical [5, 6, 7, 8] and electrical [9] excitation.
In addition to the difference in effective masses, there are several important distinctions between polaritonic and atomic systems.
First, polaritons can be directly created optically and have finite lifetimes because of the possibility for the photons to leave the system through the partially transparent Bragg mirrors.
Second, the presence of the excitonic fraction in a polariton makes possible efficient polariton-phonon interaction, which can both couple polaritons to an ensemble of incoherent excitons and lead to the energy relaxation within the polariton liquid itself.
These differences require substantial modifications of the dynamic equation 1. The finite lifetime of polaritons and an external coherent pump can also be introduced straightforwardly as a simple linear decay term and a complex time exponent source on the right-hand side of the equation 1 [10]. The incoherent pump and the coupling between polaritons and incoherent excitons are usually described in the framework of the Wouters-Carusotto model, which was first formulated for scaler polaritons [11] and then generalized for the spinor case [12].
Regarding the relaxation of energy within the condensate, there have been attempts, following the original approach of Pitaevskii [13], to include it in the phenomenological model by making the Hamiltonian non-Hermitian as $\hat{H}\rightarrow(1-\Lambda)\hat{H}$, where $\Lambda$ is a small dimensionless parameter that controls the strength of dissipation [14]. However, it can be easily seen that this energy-dependent damping does not conserve the total number of particles in the condensate and therefore cannot be viewed as pure energy relaxation.
The present paper represents an attempt to construct the theory of pure energy relaxation in a polariton system. We base our consideration on the quantum hydrodynamics approach, describing the system dynamics via set of classical field Hamilton equations for canonically conjugated variables of concentration and phase two and introducing relaxation in a natural way by adding the gradient term. We analyze how this term affects the dynamics of the polariton droplets and dispersion of the elementary excitations. We consider the scalar case, focusing on pure energy relaxation only, leaving the spinor case and a description of polarization and spin relaxation for follow-up work.
Our goal here is not to present a modeling of system dynamics in any particular experimental configuration, but rather to focus on the fundamental role of the damping mechanisms, which were overlooked before. Therefore, for the reason of clarity of the presentation, in our analysis we neglect all other dissipative processes, such as finite lifetimes, external pumping, and coupling with an incoherent excitonic reservoir, for which well established theoretical approaches exist already. These terms can be easily taken into account when analyzing a particular set of experimental data.
[…]
$\mathscr{L}=\frac{i\hbar}{2}\left(\psi^{*}\frac{\partial\psi}{\partial t}-\psi\frac{\partial\psi^{*}}{\partial t}\right)-\frac{\hbar^{2}}{2m}(\nabla\psi^{*})(\nabla\psi)-\frac{g}{2}\left(\psi^{*}\psi\right)^{2}.$ (3)
Using the Madelung representation of a field function in terms of density-phase variables,
$\psi=\sqrt{\rho}e^{-i\theta},$ (4)
one gets:
$\mathscr{L}=\hbar\rho\partial_{t}\theta-\frac{\hbar^{2}}{2m}\left[(\nabla\sqrt{\rho})^{2}+\rho(\nabla\theta)^{2}\right]-\frac{g}{2}\rho^{2}.$ (5)
From this expression it follows that the angle $\theta$ can be considered as a generalized field coordinate, while the density $\rho$ corresponds to canonically conjugated momentum as:
$\pi=\frac{\partial\mathscr{L}}{\partial(\partial_{t}\theta)}=\hbar\rho$ (6)
with the canonical field Hamiltonian being
$\mathscr{H}=\pi\partial_{t}\theta-\mathscr{L}=\frac{\hbar^{2}}{2m}\left[(\nabla\sqrt{\rho})^{2}+\rho(\nabla\theta)^{2}\right]+\frac{g}{2}\rho^{2}.$ (7)
The dynamic field equation can be thus represented in the Hamiltionan form as:
$\partial_{t}\pi=\hbar\partial_{t}\rho=-\frac{\delta\mathscr{H}}{\delta\theta}=\frac{\hbar^{2}}{m}\nabla(\rho\nabla\theta),$ (8)
$\partial_{t}\theta=\frac{\delta\mathscr{H}}{\delta\pi}=\frac{\hbar(\nabla\theta)^{2}}{2m}-\frac{\hbar}{2m\sqrt{\rho}}\nabla^{2}\sqrt{\rho}+\frac{g}{\hbar}\rho.$ (9)
This system, of course, is fully equivalent to Eq.1 (with $U=0$) and can be obtained from it directly using the substitution 4. However, it has a very important advantage as pure energy relaxation can be directly introduced to it, as it is discussed below.
Let us first note that Eq. 8 is nothing but a continuity equation for the conserving quantity $\rho$ with current density
$\mathbf{j}=-\frac{\hbar\rho}{m}\nabla\theta.$ (10)
Pure energy relaxation should not affect particle number conservation, so we can do nothing but leave this equation as is.
On the other hand, energy relaxation can be introduced in Eq. 9 via additional term proportional to $-\delta\mathscr{H}/\delta\theta$, which drags the system towards the state of the minimal energy, so that the modified dynamic equations read:
$\hbar\partial_{t}\rho=-\frac{\delta\mathscr{H}}{\delta\theta}=\frac{\hbar^{2}}{m}\nabla(\rho\nabla\theta),$ (11)
$\partial_{t}\theta=\frac{\delta\mathscr{H}}{\delta\pi}-\gamma\frac{\delta\mathscr{H}}{\delta\theta}=$ (12)
$=\frac{\hbar(\nabla\theta)^{2}}{2m}-\frac{\hbar}{2m\sqrt{\rho}}\nabla^{2}\sqrt{\rho}+\frac{g}{\hbar}\rho+\lambda\nabla(\rho\nabla\theta),$
where $\lambda=\hbar^{2}\gamma/m$ is a phenomenological energy damping constant. The energy of the system is not any more constant but decreases with time as
$\frac{dE}{dt}=\int\frac{\partial\mathscr{H}}{\partial t}d^{2}\mathbf{r}=-\frac{\lambda\hbar^{2}}{m}\int\left[\nabla(\rho\nabla\theta)\right]^{2}d^{2}\mathbf{r}=-\lambda m\int|\nabla\mathbf{j}|^{2}d^{2}\mathbf{r}\leq 0.$ (13)
The system thus relaxes to the minimal energy state, conserving the total number of particles.
Note that we can rewrite the equations 11,12 back to a single equation for the field function 4 as follows:
$i\hbar\frac{\partial\psi}{\partial t}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi+g|\psi|^{2}\psi-m\lambda\psi\nabla\mathbf{j}=-\frac{\hbar^{2}}{2m}\nabla^{2}\psi+g|\psi|^{2}\psi+\frac{i\hbar}{2}\lambda\psi\left[\psi^{\ast}\nabla^{2}\psi-\psi\nabla^{2}\psi^{\ast}\right].$ (14)
This equation constitutes the main result of the current work. In the following, we show that the energy relaxation crucially affects both the dispersion of elementary excitations and the spatio-temporal dynamics of polarition droplets.
Dispersion of elementary excitations and superfluidity. It is a known fact that the dispersion of elementary excitations of a conservative spatially homogeneous condensate with density $\rho_{0}$ is gapless and linear in $\mathbf{k}$, and at small momenta is described by the widely known Bogoliubov formula $\omega(k)=\sqrt{g\rho_{0}/m}k$, which according to the Landau criterium corresponds to the onset of superfluidity in the system, with
$v_{c}=\sqrt{\frac{g\rho_{0}}{m}}$ (15)
being a critical velocity.
This result can be obtained directly from Eq. 1 by calculating the dispersion of small excitation on the background of the spatially uniform condensate. We can apply a similar procedure to analyze how the pure energy relaxation will affect the dispersion.
The total field corresponding to a condensate of a density $\rho_{0}$ and an excitation with wavevector $\bf{q}$ and frequency $\omega$ is:
$\psi(\mathbf{r},t)=e^{-i\frac{g\rho_{0}}{\hbar}t}\left[\sqrt{\rho_{0}}+\xi e^{i(\mathbf{q}\cdot\mathbf{r}-\omega t)}+\bar{\xi}^{\ast}e^{-i(\mathbf{q}\cdot\mathbf{r}-\omega t)}\right],$ (16)
where $\xi,$ and $\bar{\xi}$ are the amplitudes of the small perturbations, $|\xi|^{2},\,|\bar{\xi}|^{2}\ll\rho_{0}$. Placing ansatz Eq.16 in Eq.14 and linearizing it, one gets the following for the dispersion of elementary excitation:
$\omega(q)=\sqrt{\frac{g\rho_{0}}{m}q^{2}+\left(\frac{\hbar^{2}}{4m^{2}}-\frac{\lambda^{2}\rho_{0}^{2}}{4}\right)q^{4}}-i\frac{\lambda\rho_{0}q^{2}}{2},$ (17)
The real and imaginary parts of the dispersions are shown in Figure 1. Note that in small $q$ the real part of the dispersion remains linear $\partial\omega/\partial q|_{q=0}=v_{c}$ and therefore, according to the Landau criterion, the condensate is superfluid with critical velocity $v_{c}$ independently of $\lambda$. The negative imaginary part is responsible for the decay of the excitations provided by the pure energy relaxation term, which scales quadratically with $q$ and is proportional to the condensate density $\rho_{0}$, which, as expected, reflects the effect of bosonic stimulation.
Let us remark that both the effect of polariton superfluidity [7, 8] and the formation of a linear polariton dispersion above the condensation threshold in a dissipative system was reported experimentally [15, 16]. Note also that models of dissipative polariton fluids without pure energy relaxation give a qualitatively different dispersion, with flat regions in energy bands [11, 17], and no clean superfluid behavior [18].
Figure 1: The real (a) and imaginary (b) parts of the dispersion $E(q)=\hbar\omega(q)$ for different condensate densities $\rho_{0}$. The condensate is resting in the laboratory reference system (the phase gradient of $\psi$ is equal to zero). The real part of dispersion shows linear scaling with $q$ at small momenta, characteristic of superfluid behavior. The critical superfluid velocity increases with the density, as expected. The imaginary part of dispersion is responsible for the decay of the excitations provided by the pure energy relaxation term and scales quadratically with $q$. Parameters of the system are $g=6\cdot 10^{-3}\ meV\cdot\mu m^{2}$, $m=5\cdot 10^{-5}m_{0}$, $\rho_{0}=100\ \mu m^{-2}$, $\lambda=0.5\cdot 10^{-4}\mu\mathrm{m^{4}}\cdot\mathrm{ps}^{-1}$.
The onset of a superfluid behavior in our model can be directly tested numerically. Consider an obstacle moving across a uniform condensate corresponding to a field function $\psi=\sqrt{\rho_{0}}$ and described by the external potential in the form of a Gaussian function $V=V_{0}\mathrm{exp}\left[-(x-v_{p}t)^{2}/w_{p}^{2}\right]$ where $V_{0}$, $w_{p}$ and $v_{p}$ are the depth, width and velocity of the potential.
The results of the 2D simulations are shown in Fig. 2. The four upper panels illustrate the supersonic cases of obstacle motion in the absence (a),(b) and in the presence $\lambda=1.4\cdot 10^{-3}\ \mu\mathrm{m}^{4}\cdot\mathrm{ps}^{-1}$ of the energy relaxation. The cones of the emitted waves are clearly seen in panels (a),(c). In the corresponding spatial spectra, see panels (b), (d), there are characteristic patterns corresponding to the phase matching condition $\mathrm{Re}[\omega(\mathbf{q})]=v_{p}q_{x}$, where $\bf{q}$ is a wavevector of a scattered wave. The difference introduced by the energy relaxation is that in this case the emitted waves slowly decay so that the field becomes localized (c), and the amplitudes of the scattered waves with high momenta are suppressed (compare panels (b) and (d)), since the decay rate of the linear excitations with $\bf q$ is proportional to $\mathrm{Im}[\omega(\bf q)]\sim q^{2}$. In the case of the subsonic regime, no emitted waves have been observed, and this regime is very similar to that known for unperturbed GPE (see panels (e) and (f)).
Figure 2: The density (a) and the spatial spectrum (b) of a supersonic condensate in the presence of an obstacle moving across it, with perturbation potential being $V=V_{0}\mathrm{exp}\left[-(x-v_{p}t)^{2}/w_{p}^{2}\right]$. The density of the condensate is $\rho_{0}=50\ \mu m^{-2}$, $V_{0}=0.3\ meV$, $w_{p}=0.56\ \mu m$, $v_{p}=2.5\ \mu\mathrm{m}\cdot\mathrm{ps}^{-1}$. The nonlinear polariton-polariton interaction $g=6\cdot 10^{-3}\ m\mathrm{eV}\cdot\mu\mathrm{m}^{2}$, the relaxation rate $\lambda=0$. Panels (c) and (d) show the same but in the presence of the relaxation $\lambda=1.4\cdot 10^{-3}\mu\mathrm{m}^{4}\cdot\mathrm{ps}^{-1}$. Panels (e) and (f) show the same as (c) and (d), but for the subsonic case with $v_{p}=0.5\ \mu\mathrm{m}\cdot\mathrm{ps}^{-1}$.
Stability of superfluid and non-superfluid flows.
All stationary solutions of GPE are also solutions of (14). This follows from the fact that for that solution $\nabla\bf{j}=0$ and the term accounting for the relaxation vanishes. However, relaxation does affect the stability of solutions. To analyze this, one should use Eq. 14 the ansatz corresponding to a spatially uniform condensate propagates along the $x$ axis with the wavevector of the absolute value $\kappa$,
$\psi(\mathbf{r},t)=e^{-i(\frac{g\rho_{0}}{\hbar}+\frac{\hbar\kappa^{2}}{2m})t+i\kappa x}\left[\sqrt{\rho_{0}}+\xi e^{i(\mathbf{q}\cdot\mathbf{r}-\omega t)}+\bar{\xi}^{\ast}e^{-i(\mathbf{q}\cdot\mathbf{r}-\omega t)}\right]$ (18)
and then perform the linearization procedure. This gives the following dispersion relation for the excitations:
$\omega(q)=\sqrt{\frac{g\rho_{0}}{m}q^{2}+\left(\frac{\hbar^{2}}{4m^{2}}-\frac{\lambda^{2}\rho_{0}^{2}}{4}\right)q^{4}-\frac{i\hbar q^{2}}{m}\lambda\rho_{0}\kappa q_{x}}-i\frac{\lambda\rho_{0}q^{2}}{2}+\frac{\hbar}{m}\kappa q_{x},$ (19)
Note that for $\kappa=0$ Eqs.17 and 19 are equivalent. However, differently from Eq.17 the imaginary part of Eq.19 can become positive, resulting in the development of the flow instability. The typical dependence of the instability increment (positive imaginary part of (17)) for an unstable condensate is shown in Fig. 3(a) as a function of $q_{x}$, $q_{y}$.
Figure 3: Panel (a): increment rate of the perturbations characterized by the wavevectors $q_{x}$ and $q_{y}$ (imaginary part of $\omega(\vec{q})$ given by (17)). The density of the condensate is $\rho_{0}=150\ \mu\mathrm{m}^{-2}$ , the condensate propagates along the $x$ axis with the wavevector $\kappa=1.3\ \mu\mathrm{m}^{-1}$. The parameters are $g=6\cdot 10^{-3}\ m\mathrm{eV}\cdot\mu\mathrm{m}^{2}$, $\lambda=1.4\cdot 10^{-3}\mu\mathrm{m^{4}}\cdot\mathrm{ps}^{-1}$. The white lines show the boundary where the increments roll to zero. Panel (b): maximum increment of the instability as a function of the condensate density $\rho_{0}$ and its wavevector $\kappa$. The white line shows the border separating the stable condensate (above the curve) and unstable condensate (below the curve). At $\rho=0$ the condensate is neutrally stable.
The dependency of the maximum instability increment as a function of $\rho_{0}$ and $\kappa$ is shown in Fig. 3(b). As expected, it is higher for fast-moving condensates. However, the dependency on the condensate density $\rho_{0}$ at fixed $\kappa$ is not monotonous: it is always zero for $\rho_{0}=0$, then grows, reaches its maximum, and then at some density is rolled to zero so that the condensate stabilizes. It happens exactly at the point where the velocity of the condensate becomes equal to the critical velocity defined by the Landau criterion. More details on the development of the instability are given in Supp. Mat.
Deceleration of polariton droplets. Let us now analyze how energy relaxation affects the dynamics of polariton droplets. To keep the presentation short, we consider the simplest case of one-dimensional Gaussian wavepacket $\psi(t=0,x)=\sqrt{\rho_{m}}\exp(-x^{2}/w_{0}^{2}+ik_{0}x)$ with the maximum density $\rho_{m}$ and wavevector $k_{0}$.
[…]
Figure 4: Panel (a): dependencies of the condensate droplet velocity $v$ on the propagation time $t$ for initial wavevector $k_{0}=0.7\ \mu\mathrm{m}^{-1}$ and different initial peak densities $\rho_{m}=50\ \mu\mathrm{m}^{-2}$ (black curve), $\rho_{m}=125\ \mu\mathrm{m}^{-2}$ (blue curve) and $\rho_{m}=200\ \mu\mathrm{m}^{-2}$ (red curve). Panel (b) shows the same, but for the initial wavevector $k_{0}=2\ \mu\mathrm{m}^{-1}$. For all cases $g=0$ and $\lambda=1.4\cdot 10^{-3}\mu\mathrm{m}^{4}\cdot\mathrm{ps}^{-1}$.
As expected, the deceleration rate increases with the condensate densities; see Supp. Mat. We have checked that at short propagation times the velocity decreased exponentially, $v=v_{0}\mathrm{exp}(-\gamma t)$ with the decay rate $\gamma$ being independent on $k_{0}$ and scaling linearly with $\rho_{m}$. At longer propagation distances the pulse broadens, its intensity decreases, and the velocity scaling with time becomes polynomial, $v\sim t^{-2}$. Polartion-polariton interactions are repelling, so that they speed up the condensate spreading and thus suppress its deceleration. A detailed discussion of these effects will be presented elsewhere.
Conclusion We have shown that the pure energy relaxation can be naturally introduced into the Gross-Pitaevskii equation. The resulting dissipation term conserves the number of particles and does not destroy the effect of superfluidity but strongly affects the dynamics of polariton droplets. Our results can also be applied to the cases of other bosonic condensates where the effects of energy relaxation are important.